OpenStudy (goformit100):

Using Euclid’s division algorithm, find which of the following pairs of numbers are co-prime: (i) 231, 396 (ii) 847, 2160

OpenStudy (goformit100):

@satellite73

OpenStudy (anonymous):

you know how to use it?

mathslover (mathslover):

Dividend = Divisor * quotient + Remainder or a = bq + r Divide 396 by 231 396 = 231 * 1 + 156 Again divide 231 by 156 (as Remainder = 156 and not equal to zero) 231 = 156 * 1 + 75 . . . Do it unless u get remainder = 0

OpenStudy (goformit100):

Thanks

OpenStudy (zarkon):

165 not 156

mathslover (mathslover):

Pardon :( @Zarkon , thanks for pointing it out.

mathslover (mathslover):

@goformit100 , do the process from start then!

OpenStudy (goformit100):

ok

OpenStudy (goformit100):

ok sure Sir

mathslover (mathslover):

If divisor is 1, then they are co prime , if divisor is not 1 , then they are not coprime. [Notice the divisor when remainder becomes zero]

OpenStudy (goformit100):

ya